# Beyond Characteristic Equations: A Unified Non-Bloch Band Theory via Wavefunction Data

### v1

@@ Abstract
Introduction.— In recent years, there has been a remarkable surge in research interest surrounding non-Hermitian systems. Such systems naturally arise in open quantum systems [1, 2, 3, 4, 5] and in photonic setups with gain and loss [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Among various non-Hermitian phenomena, the non-Hermitian skin effect (NHSE) [17, 18, 19, 20, 21, 22, 23] has attracted particular attention due to its profound implications for fundamental physical principles. In particular, the NHSE leads to a striking breakdown of conventional Bloch band theory and a collapse of the bulk-boundary correspondence, thereby challenging the foundational understanding of band structures in Hermitian systems.
@@
The breakdown of conventional Bloch band theory in predicting open-boundary energy spectra poses a fundamental challenge in non-Hermitian system studies, demanding innovative approaches to compute the energy spectra. The theory that addresses this issue is referred to as non-Bloch band theory [17, 24, 25, 26]. The main idea can be summarized as follows: replacing the conventional Bloch wavevector through the substitution $e^{ik}\rightarrow\beta$, thereby mapping the original non-Hermitian Bloch Hamiltonian $H(k)$ to its non-Bloch counterpart $H(\beta)$. The spectral determination process hinges on solving the characteristic equation $\text{det}[E-H(\beta)]=0$. By solving the zeros of this equation with respect to $\beta$ and applying the condition to get the continuum bands, one can trace out a complex manifold of $\beta$ values, which is referred to as the generalized Brillouin zone (GBZ) [17, 18, 24, 27, 28, 29, 30, 31, 32, 33, 34, 35]. This complex-plane trajectory fundamentally serving as the spectral determinant that replaces conventional Brillouin zone concepts. For convenience, we call the condition to get the continuum bands the GBZ condition. The above description tells that the input information of the non-Bloch theory is the characteristic function $\text{det}[E-H(\beta)]$, and once this is determined, the energy spectrum “seemingly” can be obtained through a standard procedure.
@@
Although the GBZ condition derived in non-Bloch band theory successfully applies to many systems [24], it has been shown to fail in certain cases where specific symmetries are present [30, 36, 37, 38]. A common remedy is to modify the GBZ condition in a manner tailored to the corresponding symmetry. However, we find this approach largely ad hoc and lacking in generality, which we regard as unsatisfactory. In parallel, other studies have explored this issue in more general contexts based on the algebraic structure of the characteristic function $\text{det}[E-H(\beta)]$ [27, 39]. These works suggest that changes in the GBZ condition can be systematically understood through this algebraic structure. In our recent investigation, we present a counterexample that challenges this interpretation: we identify systems that share the same non-Bloch characteristic equation yet exhibit distinct energy spectra and GBZs. This finding indicates that the characteristic equation alone is insufficient to fully determine the energy spectrum in certain cases, pointing to a fundamental limitation of the current non-Bloch band theory.
@@
This naturally raises the question: what additional information is required, beyond the characteristic function $\text{det}[E-H(\beta)]$, to fully determine the energy spectrum under open boundary conditions? The answer actually lies in the wavefunction data. Here, the wavefunction data means the following: the roots $\beta$ of the characteristic equation, often referred to as generalized eigenvalues, extend the notion of eigenvalues in linear algebra and encode the spatial decay or growth of the wavefunction. Correspondingly, one can define generalized eigenvectors $v(\beta)$ that satisfy $[E-H(\beta)]v(\beta)=0$, capturing the internal structure of the wavefunction within a unit cell. We find that the pair $(\beta,v(\beta))$ provides a complete description of the energy spectrum. Within this framework, we achieve a unified formulation of non-Bloch band theory that applies to systems with arbitrary symmetries. Moreover, this approach allows for the straightforward algebraic computation of quantities such as the end-to-end signal amplification factor, which is often difficult to evaluate in systems with high symmetry.
@@
Figure 1: The open boundary spectra and GBZs of two models, $H_{1}(k)$ and $H_{2}(k)$, which share the same characteristic function. The parameters are set as $t_{0}=1$ (energy unit) and $a=b=c=d=1$ (dimensionless). (a), (c), (e) correspond to properties of $H_{1}$, while (b), (d), (f) correspond to those of $H_{2}$. (a), (b) Numerically computed spectra for a finite chain of length $L=100$. The energy $E_{0}=-1/2$ lies within the spectrum of $H_{1}$ but not within that of $H_{2}$. (c), (d) The GBZs of two models, shown as blue curves. Both models share the same auxiliary GBZ, plotted as faint red curves for reference. Notably, the GBZ in (d) includes $\beta=0$. In the GBZ of $H_{1}$, the mode with generalized momentum $\beta_{0}$ corresponds to the energy $E_{0}$, whereas $\beta_{0}$ does not lie on the GBZ of $H_{2}$. (e), (f) Spectra reconstructed from the GBZs (c) and (d), respectively, representing the open boundary spectra in the limit $L\rightarrow\infty$.
Previous studies have all taken the characteristic function as their starting point, whereas our work is motivated by a counterexample demonstrating that using the characteristic function alone as the input of the theory is fundamentally incomplete. Fig. 1 shows that two triple bands non-Hermitian Hamiltonians with the same characteristic function. Fig. 1(a) is the open boundary spectrum of the Bloch Hamiltonian $H_{1}(k)=\left(\begin{array}[]{ccc}t_{0}e^{ik}&b\,t_{0}&0\\ a\,t_{0}e^{ik}&t_{0}e^{ik}&d\,t_{0}e^{-ik}\\ 0&c\,t_{0}e^{-ik}&t_{0}e^{ik}\\ \end{array}\right)$ and Fig. 1(b) is the open boundary spectrum of the Bloch Hamiltonian $H_{2}(k)=\left(\begin{array}[]{ccc}t_{0}e^{ik}&0&0\\ 0&t_{0}e^{ik}&t_{0}\left(a\,b\,e^{2ik}+c\,d\,e^{-ik}\right)\\ 0&t_{0}e^{-ik}&t_{0}e^{ik}\\ \end{array}\right)$. By performing a variable substitution $e^{ik}\rightarrow\beta$, both models result in the same characteristic function $\text{det}[E-H_{i}(\beta)]=\left(E-t_{0}\beta\right)\left(t_{0}^{2}\beta^{2}-2t_{0}E\beta-abt_{0}^{2}\beta+E^{2}-\frac{cdt_{0}^{2}}{\beta^{2}}\right),i=1,2$. However, the energy spectra of the two are not only different but also cannot be contained within each other. This difference highlights a fundamental limitation of non-Bloch band theory when the characteristic function is taken as the sole input: the theory is incomplete. One may observe that $H_{2}$ is a decoupled Hamiltonian, which causes its GBZ condition differs from the conventional form. While this observation is correct, a more accurate interpretation is that the symmetry of $H_{2}$ enforces a modification of the GBZ condition. Indeed, it is not only decoupling but also other types of symmetry, such as the symplectic symmetry, that can alter the GBZ condition.
@@
Review of non-Bloch band theory.— Start from a one-dimensional tight-binding model in an open chain. Assume its Bloch Hamiltonian is given by $H(k)=\sum_{n=-m}^{m}a_{n}e^{ikn}$ with $a_{-m},a_{-m+1},\cdots,a_{m}$ being $12m+$ 1 hopping matrices with dimension $l$, and the corresponding tight-binding Hamiltonian is $H=\sum_{l}\sum_{n=-m}^{m}a_{n}c_{l+n}^{\dagger}c_{l}$. The ansatz for the wavefunction is assumed to be $\psi(n)=\beta^{-n}\psi(0)$, where $\psi(n)$ represents the wavefunction amplitude at the $n$-th unit cell and $\beta$ represents the spatial decay rate. Due to the translation symmetry in the bulk, the bulk equation is a characteristic equation
@@
where $H(\beta)$ is a reformulation of $H(k)$ by performing a variable substitution $e^{ik}\rightarrow\beta$. This is an algebraic equation of both $\beta$ and $E$. For each $E$, there are $2M$ ($M=ml$) roots of $\beta$ by taking account the zero root and the root at infinity, which can be sorted by their magnitudes: $|\beta_{1}(E)|\leqslant|\beta_{2}(E)|\leqslant\cdots\leqslant|\beta_{2M}(E)|$. The boundary condition gives extra equations, which involve the following [24]
@@
where the set $P$ is a subset of the mode index set $\{1,2,\cdots,2M\}$ with $M$ elements, $F_{P}$ is a scalar for each $P$, and $L$ is the length of the one-dimensional chain. To produce the continuum band, the magnitude of the leading-order term and the next-to-leading-order terms of Eq. (2) must be equal. This requirement leads to the condition $|\beta_{2M}\beta_{2M-1}\cdots\beta_{M+2}\beta_{M+1}|=|\beta_{2M}\beta_{2M-1}\cdots\beta_{M+2}\beta_{M}|$, which is equivalent to $|\beta_{M}|=|\beta_{M+1}|$ [24]. This condition is known as the GBZ condition, as it determines the shape of the GBZ.
@@
As we have shown in Fig. 1, changes in the GBZ condition of this kind cannot be inferred from the characteristic function $\text{det}[E-H(\beta)]$. In other word, the vanishing of the coefficients $F_{P}$ is not encoded in the characteristic function, and the central contributions of this work is a criterion for determining when $F_{P}=0$ by using the extra information beyond the characteristic function, which will be introduced in the following.
@@
In other word, the wavefunction ansatz of the mode $i$ is $\psi(n)=\beta_{i}^{-n}v_{i}$. It is quite easy to solve $v_{i}$ since $[E-H(\beta_{i})]$ is a matrix with dimension $l$, which is a small size. For convenience, we call $\beta_{i}$ the generalized eigenvalues of the characteristic equation and $v_{i}$ the generalized eigenvector of the generalized eigenvalue $\beta_{i}$. Similarly, there are also left generalized eigenvectors $u_{i}$ satisfy
@@
Left/Right generalized eigenvectors all encode the wavefunction information. In the following, we explain how to express the GBZ condition in terms of these wavefunction information. First, note that $M=ml$ where $l$ the dimension of Hamiltonian $H(\beta)$. We can construct supercell right generalized eigenvectors
@@
It can be seen that they are corresponding wavefunctions inside the supercell with $m$ unit cells, which is quite easy to calculate since $m=M/l$ is quite small. The above form relys on the assumption that all $\beta_{i}$ are nondegenerate, and the form of the degenerate case can be seen in the Supplemental Material (SM) [40]. Now we can express the criterion for determining when $F_{P}=0$ in Eq. (2), i.e.,
@@
There are $M$ terms in $\sum_{i\in P}x_{i}y_{i}^{T}$ and each term $x_{i}y_{i}^{T}$ is a $M\times M$ matrix of rank 1. Hence Eq. (7) can be understood as a biorthogonal completeness relation, it tells whether or not these $M$ modes from the subset $P$ can span the whole Hilbert space. Our additional GBZ condition states that for the term $\prod_{i\in P}(\beta_{i})^{L}$ be the leading term of Eq. (2), the corresponding modes from index set $P$ must form a complete basis, which is expressed by the mathematical equation $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)\neq 0$. In other word, the leading term of Eq. (2) is the one for which $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)\neq 0$ and the product $\prod_{k\in P}|\beta_{k}|$ is maximized, which can be mathematically expressed as
@@
where $P_{\text{max}}$ denotes the set of mode indices for the leading term, while “$\mathop{\mathrm{arg}\,\max}$” represents the standard mathematical operator for identifying indices of maximal elements - a notation widely used in optimization, mathematics, and computer science. From Eq. (8), it can be seen that Eq. (7) is the key result of this paper, and we put the mathematical derivation in SM [40].
@@
Based on Eq. (8), the algorithm of the modified non-Bloch band theory can be summarized as Fig. 2(b), and can be compared to the original non-Bloch band theory shown in Fig. 2(a).
@@
Figure 2: (a) Workflow of the original non-Bloch band theory algorithm. Only the most common scenario is shown; in systems with symmetry, the GBZ condition may require case-specific treatment. (b) Workflow of the modified non-Bloch band theory algorithm. The application of this algorithm to specific models $H_{1}$ and $H_{2}$ is illustrated in Table 1 and Table 2.
@@
The above theoretical framework may seem abstract, we now validate it using the previously discussed triple-band models, as shown in Fig. 1. In Fig. 1(a), we demonstrate that a specific energy level, $E_{0}=-1/2$, lies within the spectrum of $H_{1}$, whereas Fig. 1(b) shows that it does not belong to the spectrum of $H_{2}$. This discrepancy can be explained using the completeness condition in Eq. (7). By substituting $E=-1/2$ into the characteristic equation Eq. (1), we obtain 6 roots: $\beta_{1}=0,\beta_{2}=-0.5,\beta_{3}=-0.94,\beta_{4}=0.94,\beta_{5}=-1.06i,\beta_{6}=1.06i$, ordered by their magnitudes. For each three-element subset $P$ of the index set $\{1,2,\cdots,6\}$, we test the criterion in Eq. (7). In Table 1 and Table 2, we list several such index subsets $P$ that yield large values of $\prod_{i\in P}|\beta_{i}|$, corresponding to Hamiltonians $H_{1}$ and $H_{2}$, respectively. According to Eq. (8), the leading-order term and the next-to-leading-order term are highlighted in red. From Table 1, it can be seen that $\beta_{4}\beta_{5}\beta_{6}$ and $\beta_{3}\beta_{5}\beta_{6}$ serve as the leading and next-to-leading-order terms, and the GBZ condition $|\beta_{4}\beta_{5}\beta_{6}|=|\beta_{3}\beta_{5}\beta_{6}|$ is satisfied. This indicates that $E=-1/2$ lies in the spectrum of $H_{1}$. In contrast, Table 2 shows that for $H_{2}$, the leading and next-to-leading-order terms are given by $\beta_{2}\beta_{5}\beta_{6}$ and $\beta_{2}\beta_{4}\beta_{6}$, respectively. However, the GBZ condition $|\beta_{2}\beta_{5}\beta_{6}|=|\beta_{2}\beta_{4}\beta_{6}|$ is not satisfied, which implies that $E=-1/2$ is not part of the spectrum of $H_{2}$. Therefore, Eq. (7) is crucial to determine the GBZ condition and further the GBZ. Based on this criterion, we compute and plot the GBZs for $H_{1}$ and $H_{2}$ in Fig. 1(c) and Fig. 1(d), respectively. The spectra made from these two GBZs which represent the open boundary spectra at the chain length limit $L\rightarrow\infty$ are shown in Fig. 1(e) and Fig. 1(f).
@@
rank$\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ 3 3 3 3 3 3
det$\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$
Table 1: Biorthogonal completeness of 3 modes of $H_{1}$ at energy $E=-1/2$. The index sets $P$ of the leading-order term and the next-to-leading-order terms are highlighted in red. The energy $E=-1/2$ lies within the spectrum since the magnitude of the leading-order term and the next-to-leading-order terms are equal.
@@
rank$\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ 2 2 2 2 3 3
det$\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ $=0$ $=0$ $=0$ $=0$ $\neq 0$ $\neq 0$
Table 2: Biorthogonal completeness of 3 modes of $H_{2}$ at energy $E=-1/2$. The index set $P$ of the leading-order term and the next-to-leading-order terms are highlighted in red. The energy $E=-1/2$ is not part of the spectrum since the magnitude of the leading-order term and the next-to-leading-order terms are not equal.
@@
[…]
@@
This result admits a straightforward interpretation: the Green’s function $G_{1L}(\omega)$ receives contributions from the modes in the index set $P_{\text{min}}$, and can be written as $G_{1L}(\omega)=\sum_{i\in P}\beta_{i}^{L}G_{i}(\omega)$. Therefore, in the large $L$ limit, the matrix norm $\|G_{1L}(\omega)\|$ is dominated by the exponential growth associated with the largest $\beta_{i}$ among these modes, i.e., $\alpha_{\leftarrow}=\mathop{\max}_{i\in P_{\text{min}}}|\beta_{i}|$.

### v2

# Beyond characteristic equations: A unified one-dimensional non-Bloch band theory via wavefunction data

@@ Abstract
Introduction.— In recent years, there has been a remarkable surge in research interest surrounding non-Hermitian systems. Such systems naturally arise in open quantum systems [1, 2, 3, 4, 5] and in photonic setups with gain and loss [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The distinctive spectral characteristics of non-Hermitian systems, particularly the emergence of exceptional points, have been recognized as playing a central role in a wide range of applications including lasing and coherent perfect absorption [17, 18, 19, 20, 21]. Among various non-Hermitian phenomena, the non-Hermitian skin effect (NHSE) [22, 23, 24, 25, 26, 27, 28] has attracted particular attention due to its profound implications for fundamental physical principles. In particular, the NHSE leads to a striking breakdown of conventional Bloch band theory and a collapse of the bulk-boundary correspondence, thereby challenging the foundational understanding of band structures in Hermitian systems.
@@
The breakdown of conventional Bloch band theory in predicting open-boundary energy spectra poses a fundamental challenge in non-Hermitian system studies, demanding innovative approaches to compute the energy spectra. The theory that addresses this issue is referred to as non-Bloch band theory [22, 29, 30, 31]. The main idea can be summarized as follows: replace the conventional Bloch wavevector through the substitution $e^{ik}\rightarrow\beta$, thereby mapping the original non-Hermitian Bloch Hamiltonian $H(k)$ to its non-Bloch counterpart $H(\beta)$. The spectral determination process hinges on solving the characteristic equation $\text{det}[E-H(\beta)]=0$. By solving the zeros of this equation with respect to $\beta$ and applying the condition to get the continuum bands, one can trace out a complex manifold of $\beta$ values, which is referred to as the generalized Brillouin zone (GBZ) [22, 23, 29, 32, 33, 34, 35, 36, 37, 38, 39, 40]. This complex-plane trajectory fundamentally serves as the spectral determinant that replaces conventional Brillouin zone concepts. For convenience, we call the condition to get the continuum bands GBZ condition. The above description tells that the input information of the non-Bloch theory is the characteristic function $\text{det}[E-H(\beta)]$, and once this is determined, the energy spectrum “seemingly” can be obtained through a standard procedure.
@@
Although the GBZ condition derived in non-Bloch band theory successfully applies to many systems [29], it has been shown to fail in certain cases where specific symmetries are present [35, 41, 42, 43]. A common remedy is to modify the GBZ condition in a manner tailored to the corresponding symmetry. However, we find this approach largely ad hoc and lacking in generality, which we regard as unsatisfactory. In parallel, other studies have explored this issue in more general contexts based on the algebraic structure of the characteristic function $\text{det}[E-H(\beta)]$ [32, 44]. These works suggest that changes in the GBZ condition can be systematically understood through this algebraic structure. In our recent investigation, we present a counterexample that challenges this interpretation: We identify systems that share the same non-Bloch characteristic equation, yet exhibit distinct energy spectra and GBZs. This finding indicates that the characteristic equation alone is insufficient to fully determine the energy spectrum in certain cases, pointing to a fundamental limitation of the current non-Bloch band theory.
@@
This naturally raises the question: What additional information is required, beyond the characteristic function $\text{det}[E-H(\beta)]$, to fully determine the energy spectrum under open boundary conditions? The answer actually lies in the wavefunction data. Here, the wavefunction data mean the following: The roots $\beta$ of the characteristic equation, often referred to as generalized eigenvalues, extend the notion of eigenvalues in linear algebra and encode the spatial decay or growth of the wavefunction. Correspondingly, one can define generalized eigenvectors $v(\beta)$ that satisfy $E-H(\beta)]v(\beta)=0$, capturing the internal structure of the wavefunction within a unit cell. We find that the pair $(\beta,v(\beta))$ provides a complete description of the energy spectrum. Within this framework, we achieve a unified formulation of non-Bloch band theory that applies to systems with arbitrary symmetries. Moreover, this approach allows for the straightforward algebraic computation of quantities such as the end-to-end signal amplification factor [45, 46, 47, 48, 49], which is often difficult to evaluate in systems with high symmetry. For example, in non-Hermitian systems belonging to the symplectic class, the formula for the end-to-end signal amplification factor given in Refs. [45, 46] does not hold. To the best of our knowledge, no analogous formula applicable to such systems has been reported in the literature, and in this work we provide the corresponding expression.
@@
Figure 1: The open boundary spectra and GBZs of two models, $H_{1}(k)$ and $H_{2}(k)$, which share the same characteristic function. The parameters are set as $t_{0}=1$ (energy unit) and $a=b=c=d=1$ (dimensionless). Panels (a), (c), and (e) correspond to properties of $H_{1}$, while panels (b), (d), and (f) correspond to those of $H_{2}$. (a), (b) Numerically computed spectra for a finite chain of length $L=100$. The energy $E_{0}=-1/2$ lies within the spectrum of $H_{1}$ but not within that of $H_{2}$. Faint red vertical lines represent Re($E$) = -1/2 in both panels. (c), (d) The GBZs of two models, shown as blue curves. Both models share the same auxiliary GBZ, plotted as faint red curves for reference. Notably, the GBZ in panel (d) includes $\beta=0$. In the GBZ of $H_{1}$, the mode with generalized momentum $\beta_{0}$ corresponds to the energy $E_{0}$, whereas $\beta_{0}$ does not lie on the GBZ of $H_{2}$. (e), (f) Spectra reconstructed from the GBZs (c) and (d), respectively, representing the open boundary spectra in the limit $L\rightarrow\infty$.
Previous studies have all taken the characteristic function as their starting point, whereas our work is motivated by a counterexample demonstrating that using the characteristic function alone as the input of the theory is fundamentally incomplete. Figure 1 shows two triple-band non-Hermitian Hamiltonians with the same characteristic function. Figure 1(a) is the open boundary spectrum of the Bloch Hamiltonian $H_{1}(k)=\left(\begin{array}[]{ccc}t_{0}e^{ik}&b\,t_{0}&0\\ a\,t_{0}e^{ik}&t_{0}e^{ik}&d\,t_{0}e^{-ik}\\ 0&c\,t_{0}e^{-ik}&t_{0}e^{ik}\\ \end{array}\right)$ and Fig. 1(b) is the open boundary spectrum of the Bloch Hamiltonian $H_{2}(k)=\left(\begin{array}[]{ccc}t_{0}e^{ik}&0&0\\ 0&t_{0}e^{ik}&t_{0}\left(a\,b\,e^{2ik}+c\,d\,e^{-ik}\right)\\ 0&t_{0}e^{-ik}&t_{0}e^{ik}\\ \end{array}\right)$. By performing a variable substitution $e^{ik}\rightarrow\beta$, both models result in the same characteristic function $\text{det}[E-H_{i}(\beta)]=\left(E-t_{0}\beta\right)\left(t_{0}^{2}\beta^{2}-2t_{0}E\beta-abt_{0}^{2}\beta+E^{2}-\frac{cdt_{0}^{2}}{\beta^{2}}\right),i=1,2$. However, the energy spectra of the two not only are different but also cannot be contained within each other. This difference highlights a fundamental limitation of non-Bloch band theory when the characteristic function is taken as the sole input: The theory is incomplete. One may observe that $H_{2}$ is a decoupled Hamiltonian, which causes its GBZ condition to differ from the conventional form. While this observation is correct, a more accurate interpretation is that the symmetry of $H_{2}$ enforces a modification of the GBZ condition. Indeed, it is not only decoupling but also other types of symmetry, such as the symplectic symmetry, that can alter the GBZ condition.
@@
Review of non-Bloch band theory.— Start from a one-dimensional tight-binding model in an open chain. Assume its Bloch Hamiltonian is given by $H(k)=\sum_{n=-m}^{m}a_{n}e^{ikn}$ with $a_{-m},a_{-m+1},\cdots,a_{m}$ being $2m+1$ hopping matrices with dimension $l$, and the corresponding tight-binding Hamiltonian is $H=\sum_{l}\sum_{n=-m}^{m}a_{n}c_{l+n}^{\dagger}c_{l}$. The ansatz for the wavefunction is assumed to be $\psi(n)=\beta^{-n}\psi(0)$, where $\psi(n)$ represents the wavefunction amplitude at the $n$-th unit cell and $\beta$ represents the spatial decay rate. Due to the translation symmetry in the bulk, the bulk equation is a characteristic equation
@@
where $H(\beta)$ is a reformulation of $H(k)$ by performing a variable substitution $e^{ik}\rightarrow\beta$. This is an algebraic equation of both $\beta$ and $E$. For each $E$, there are $2M$ ($M=ml$) roots of $\beta$ by taking into account the zero root and the root at infinity, which can be sorted by their magnitudes: $|\beta_{1}(E)|\leqslant|\beta_{2}(E)|\leqslant\cdots\leqslant|\beta_{2M}(E)|$. The boundary condition gives extra equations, which involve the following [29]
@@
where the set $P$ is a subset of the mode index set $\{1,2,\cdots,2M\}$ with $M$ elements, $F_{P}$ is a scalar for each $P$, and $L$ is the length of the one-dimensional chain. To produce the continuum band, the magnitude of the leading-order term and the next-to-leading-order term of Eq. (2) must be equal. This requirement leads to the condition $|\beta_{2M}\beta_{2M-1}\cdots\beta_{M+2}\beta_{M+1}|=|\beta_{2M}\beta_{2M-1}\cdots\beta_{M+2}\beta_{M}|$, which is equivalent to $|\beta_{M}|=|\beta_{M+1}|$ [29]. This condition is known as the GBZ condition, as it determines the shape of the GBZ.
@@
As we have shown in Fig. 1, changes in the GBZ condition of this kind cannot be inferred from the characteristic function $\text{det}[E-H(\beta)]$. In other words, the vanishing of the coefficients $F_{P}$ is not encoded in the characteristic function, and the central contribution of this work is a criterion for determining when $F_{P}=0$ by using the extra information beyond the characteristic function, which will be introduced in the following.
@@
In other word, the wavefunction ansatz of the mode $i$ is $\psi(n)=\beta_{i}^{-n}v_{i}$. It is quite easy to solve $v_{i}$ since $E-H(\beta_{i})]$ is a matrix with dimension $l$, which is a small size. For convenience, we call $\beta_{i}$ the generalized eigenvalues of the characteristic equation and $v_{i}$ the generalized eigenvector of the generalized eigenvalue $\beta_{i}$. Similarly, there are also left generalized eigenvectors $u_{i}$ that satisfy
@@
Left and right generalized eigenvectors all encode the wavefunction information. In the following, we explain how to express the GBZ condition in terms of this wavefunction information. First, note that $M=ml$, where $l$ is the dimension of Hamiltonian $H(\beta)$. We can construct supercell right generalized eigenvectors
@@
It can be seen that they are corresponding wavefunctions inside the supercell with $m$ unit cells, which is quite easy to calculate since $m=M/l$ is quite small. The above form relies on the assumption that all $\beta_{i}$ are nondegenerate, and the form of the degenerate case can be seen in the Supplemental Material [50], which is inspired by Ref. [51] and derived independently in Refs. [52, 53]. Now we can express the criterion for determining when $F_{P}=0$ in Eq. (2), i.e.,
@@
There are $M$ terms in $\sum_{i\in P}x_{i}y_{i}^{T}$ and each term $x_{i}y_{i}^{T}$ is a $M\times M$ matrix of rank 1. Hence Eq. (7) can be understood as a biorthogonal completeness relation, it tells whether or not these $M$ modes from the subset $P$ can span the whole Hilbert space. Our additional GBZ condition states that for the term $\prod_{i\in P}(\beta_{i})^{L}$ to be the leading term of Eq. (2), the corresponding modes from index set $P$ must form a complete basis, which is expressed by the mathematical equation $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)\neq 0$. In other words, the leading term of Eq. (2) is the one for which $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)\neq 0$ and the product $\prod_{k\in P}|\beta_{k}|$ is maximized, which can be mathematically expressed as
@@
where $P_{\text{max}}$ denotes the set of mode indices for the leading term, while $\mathop{\mathrm{arg}\,\max}$” represents the standard mathematical operator for identifying indices of maximal elements - a notation widely used in optimization, mathematics, and computer science. From Eq. (8), it can be seen that Eq. (7) is the key result of this paper, and we put the mathematical derivation in the Supplemental Material [50].
@@
Based on Eq. (8), the algorithm of the modified non-Bloch band theory can be summarized as Fig. 2(b) and can be compared to the original non-Bloch band theory shown in Fig. 2(a).
@@
Figure 2: (a) Workflow of the original non-Bloch band theory algorithm. Only the most common scenario is shown; in systems with symmetry, the GBZ condition may require case-specific treatment. (b) Workflow of the modified non-Bloch band theory algorithm. The application of this algorithm to specific models $H_{1}$ and $H_{2}$ is illustrated in Tables 1 and 2.
@@
The above theoretical framework may seem abstract; we now validate it using the previously discussed triple-band models, as shown in Fig. 1. In Fig. 1(a), we demonstrate that a specific energy level, $E_{0}=-1/2$, lies within the spectrum of $H_{1}$, whereas Fig. 1(b) shows that it does not belong to the spectrum of $H_{2}$. This discrepancy can be explained using the completeness condition in Eq. (7). By substituting E=1/2E=-1/2 into the characteristic equation [Eq. (1)], we obtain six roots: $\beta_{1}=0,\beta_{2}=-0.5,\beta_{3}=-0.94,\beta_{4}=0.94,\beta_{5}=-1.06i,\beta_{6}=1.06i$, ordered by their magnitudes. For each three-element subset $P$ of the index set $\{1,2,\cdots,6\}$, we test the criterion in Eq. (7). In Tables 1 and 2, we list several such index subsets $P$ that yield large values of $\prod_{i\in P}|\beta_{i}|$, corresponding to Hamiltonians $H_{1}$ and $H_{2}$, respectively. According to Eq. (8), the leading-order term and the next-to-leading-order term are indicated in bold. From Table 1, it can be seen that $\beta_{4}\beta_{5}\beta_{6}$ and $\beta_{3}\beta_{5}\beta_{6}$ serve as the leading-order term and next-to-leading-order term, and the GBZ condition $|\beta_{4}\beta_{5}\beta_{6}|=|\beta_{3}\beta_{5}\beta_{6}|$ is satisfied. This indicates that E=1/2E=-1/2 lies in the spectrum of $H_{1}$. In contrast, Table 2 shows that for $H_{2}$, the leading-order term and next-to-leading-order term are given by $\beta_{2}\beta_{5}\beta_{6}$ and $\beta_{2}\beta_{4}\beta_{6}$, respectively. However, the GBZ condition $|\beta_{2}\beta_{5}\beta_{6}|=|\beta_{2}\beta_{4}\beta_{6}|$ is not satisfied, which implies that E=1/2E=-1/2 is not part of the spectrum of $H_{2}$. Therefore, Eq. (7) is crucial to determine the GBZ condition and further the GBZ. Based on this criterion, we compute and plot the GBZs for $H_{1}$ and $H_{2}$ in Figs. 1(c) and 1(d), respectively. The spectra made from these two GBZs that represent the open boundary spectra at the chain length limit $L\rightarrow\infty$ are shown in Figs. 1(e) and 1(f).
@@
[…]
Table 1: Biorthogonal completeness of three modes of $H_{1}$ at energy E=1/2E=-1/2. The index sets $P$ of the leading-order term and the next-to-leading-order term are indicated in bold. The energy E=1/2E=-1/2 lies within the spectrum since the magnitudes of the leading-order term and the next-to-leading-order terms are equal.
@@
[…]
Table 2: Biorthogonal completeness of three modes of $H_{2}$ at energy E=1/2E=-1/2. The index sets $P$ of the leading-order term and the next-to-leading-order term are indicated in bold. The energy E=1/2E=-1/2 is not part of the spectrum since the magnitudes of the leading-order term and the next-to-leading-order terms are not equal.
@@
Application in end-to-end signal response.— Above we have discussed the importance of the wavefunction information in determining the energy spectrum. We now turn to discuss its role in computing the physical response, i.e., the Green’s function. The non-Hermitian Green’s function has numerous applications in physics. For instance, it captures the end-to-end response of field coherence in bosonic open systems [45]. Even in Hermitian mesoscopic transport problems, it serves as an effective mathematical tool that enables convenient computation of nonreciprocal transport [54, 55]. These problems pertain to the end-to-end response of the system. It is known that in general non-Hermitian systems, a signal injected at one end of a chain can be amplified at the other end [56]. This amplification follows an exponential law, described by the Green’s function as $\|G_{1L}(\omega)\|=\alpha_{\leftarrow}^{L}$, where $L$ denotes the chain length, $G_{1L}(\omega)=(\omega-H)^{-1}_{1L}$ is the Green’s function from the right end to the left end, and $\|\cdot\|$ represents the absolute value in single-band systems and the matrix norm in multiband systems. The exponential factor $\alpha_{\leftarrow}$ is actually $|\beta_{M}(\omega)|$ (with respect to the convention adopted here) in a general single-band non-Hermitian systems [45, 46]. Here, $|\beta_{i}(\omega)|$ are solutions of $\text{det}[\omega-H(\beta)]=0$ and $|\beta_{1}(\omega)|\leqslant|\beta_{2}(\omega)|\leqslant\cdots\leqslant|\beta_{2M}(\omega)|$. Thus, if $|\beta_{M}(\omega)|>1$, the signal is amplified from right to left. Although single-band systems are well understood, a corresponding result for general multiband systems remains lacking. For example, for a Hamiltonian belonging to the symplectic class, should $\alpha_{\leftarrow}$ be given by $|\beta_{M}(\omega)|$ or $|\beta_{M+1}(\omega)|$? To address this question, we investigate the multiband non-Hermitian Green’s function and derive a universal expression for $\alpha_{\leftarrow}$. The first step is to identify the leading contribution that satisfies the completeness relation in Eq. (7); specifically, one must determine an index subset $P_{\text{min}}$ containing $M$ modes that fulfills
@@
This result admits a straightforward interpretation: The Green’s function $G_{1L}(\omega)$ receives contributions from the modes in the index set $P_{\text{min}}$, and can be written as $G_{1L}(\omega)=\sum_{i\in P_{\text{min}}}\beta_{i}^{L}G_{i}(\omega)$. Therefore, in the large $L$ limit, the matrix norm $\|G_{1L}(\omega)\|$ is dominated by the exponential growth associated with the largest $\beta_{i}$ among these modes, i.e., $\alpha_{\leftarrow}=\mathop{\max}_{i\in P_{\text{min}}}|\beta_{i}|$.